# Algorithmic and High-Frequency Trading Problems for Semi-Markov and Hawkes Jump-Diffusion Models

### v1

@@ 4.2 Liquidation Problem Solution
For the liquidation problem, as for the acquisition problem in section 4.1, we must ensure that we have a well-posed problem. In order to do this, we now specify a boundary condition along $S_{max}\gg S_{min}$, similarly to how a boundary condition was specified along $S_{min}\ll S_{max}$ in section 4.1. Here, we specify the boundary condition $\partial_{ss}h|_{S=S_{max}}=0$. Then, the numerical scheme, again for $i=1,2,...,M-1$ and $n=0,1,...,N-1$, for the liquidation problem, where its PDE was defined in equation (43), can be derived as follows:
@@
where $M_{1}$ is defined as in section 4.1. Note that here one must also account for the boundary conditions at $i=0$, $i=M$ and for the terminal condition at $n=N$. These are as defined in equation (44)-(45) and the condition above where a upper boundary condition was set at $S_{max}$. One can see that equation (54) is very similar to equation (51), which was used to solve the acquisition problem. The main changes here are in the boundary conditions. The liquidation problem now has a price floor rather than a price cap, as in the acquisition problem.
Next, in Figure 2, we show a set of subplots which display contour plots portraying how the solution changes as a function of time and the financial asset’s price, over different values of $\bar{\sigma}$ and $\varsigma$. Time, here, is again from $t=0$ to the maturity time $t=T=1$, and the midprice (the space variable) is from $S_{min}=19.8$ to $S_{0}=20$. We specifically focus on this part of the solution as this is where it varies the most. Right of each subplot shows the color scheme for the values of the solution that are displayed. These solution values vary from 00 to $0.01$, where $\alpha=0.01$ represents the value of the terminal condition and the boundary condition at $T$ and $S_{min}$, respectively, as defined in equations (44)-(45). We can see that the agents speed of liquidating units increases as we get closer to the terminal time. This is because the agent is trying to avoid the terminal penalty cost that they must pay to liquidate any remaining units at maturity. Furthermore, the speed of trading also increases as the asset’s price approaches the floor price, $S_{min}$, again to avoid the other terminal penalty. This trading process was formalized in the stopping time in equation (35).
@@
As in Figure 1, the format for the pricing models in these plots is similar. In the top left subplot of Figure 2, we show the solution to the liquidation problem for the case where both $\bar{\sigma}=0$ and $\varsigma=0$, which would represent the case for the benchmark models in Cartea et al., (2015). In the top right subplot, we show the solution for the case where $\bar{\sigma}=0.1$ and $\varsigma=0$, which would portray the same solution were the pricing model to follow the price process in Cartea et al., (2015) combined with the price processes developed in Roldan Contreras and Swishchuk, (2022) and Roldan Contreras, (2023). The bottom two subplots, where we increased the value of the parameter $\varsigma$, then show the solutions for the cases where the price process follows our more general jump-diffusion model given in equation (17).
@@
In each consecutive subplot, from left to right, we again vary either $\bar{\sigma}$ or $\varsigma$ to show the effect that the Semi-Markov and Hawkes process parts can have on the optimal solution. In each of these subplots, one can see the effect on the solution matrix as we slowly incremented the effect of either of the parameters, $\bar{\sigma}$ or $\varsigma$, by 0.1. We can see that as they increase, the speed of trading increases at earlier time steps at lower asset prices. This makes sense, intuitively, as the higher these coefficients become, the more overall volatility the solution predicts. Thus, in these scenarios, it would then be more optimal to liquidate units of the asset quicker based on our model. In this particular type of trading problem, where the agent is liquidating units, it’s mainly price jumps that go down which can significantly change the optimal solution as the agent will want to liquidate units of the asset a lot quicker then. In the case where the asset price would jump up, this would be highly favourable to the agent, and the optimal speed of liquidating any remaining units of the asset would then mostly depend on time.
@@
Most of the input parameters for this model are kept the same as in Table 1 in section 4.1 for the acquisition problem. The only thing that we changed was $S_{min}$ and $S_{max}$, which were here set to $19.8$ and $22$, respectively. These were changed so that we could focus more on the impact of the price floor on the optimal solution, whereas in section 4.1 the solution is more focused on the effect of the price cap.
@@ 5 Strategy Simulations: Acquisition and Liquidation
In this section, we simulate the performance of the above acquisition and liquidation problems over $10,000$ different price paths, which is the total number of simulations we ran. In section 4, the optimal solution takes a static view of the markets, whereas in reality, market conditions are constantly evolving. This should be reflected in the agent’s strategy and is throughout our strategy simulations, where the strategy is updated at each time step to reflect this. Here, we will, again, specifically focus on how an increasing $\bar{\sigma}$ or $\varsigma$ can significantly change how the acquisition/liquidation strategy would evolve and how certain processes we defined in section 3 are effected. In order to setup these simulations, one must discretize the continuous-time processes introduced in section 3. In other words, the continuous-time processes in equations (14), (17), (18), (19), (21), (32)-(34) and (46)-(48) are discretized. Since both the acquisition and liquidation trading problems have already been solved numerically, as shown in section 4, this part is already given to us in a discrete form, and we can use the values in these solution matrices to guide our trading decisions in the strategy simulation. As we increment forward in time discretely, each simulation looks at the current price at every time step and then selects the matching value from the h solution matrix, which guides the agent on what the new optimal trading speed is. After a trade is made, the cash, inventory and execution price processes are updated. If either the terminal or boundary condition is breached, the simulation ends and the agent pays the terminal penalty to acquire/liquidate any remaining units of the asset. In this section we will again present subplots in the same format as in Figure 1 and 2. Recall there, that the top left subplot is for the Cartea et al., (2015) case, the top right subplot is for the Cartea et al., (2015) case combined with the cases in Roldan Contreras and Swishchuk, (2022) and Roldan Contreras, (2023), and the bottom two subplots are for our new more general price processes as defined in equation (17).
@@ 5.1 Acquisition Problem Strategy Simulations
In this section we will discuss the strategy simulations for the acquisition problem. These are very similar to the simulations performed in Cartea et al., (2015) under their pure diffusion models. First, see below, in Figure 3, the first five sample price paths from our simulations and, in Figure 4, a histogram showing the frequency of the average traded price in each of the $10,000$ simulations. In Figure 3, these 5 random price paths are based on our price process defined earlier in equation (17), where we repeat the simulations for different values of $\bar{\sigma}$ or $\varsigma$. In Figure 4, one can see how the different parameter values can influence the asset prices the agent ends up acquiring units at in our strategy simulations. The dashed black line in both Figures also shows the price cap, $S_{max}$, which if breached ends the simulation.
It is quite obvious that as we increment $\bar{\sigma}$ or $\varsigma$, prices are more likely to exhibit higher volatility, as this is still a diffusion model, but we will see next how this higher volatility coming from the jump part (via diffusion approximation) can significantly alter how the simulated trading strategy will evolve. One thing to notice from these plots is that when our boundary condition $h(t,S_{max}=20.2)$ is breached. Here, we encounter one of the times where the agent purchases the remaining units instantly, pays the terminal penalty $\alpha=0.01$ and stops trading. The other time this would occur at is when $t=T=1$. We will notice in the next set of plots how the strategy is then significantly altered each time this occurs and we will portray how this can significantly effect the overall performance of the strategy in these scenarios.
Next, in Figures 5 and 6, we show how the agents inventory and trading speed processes evolve over the five sample price paths we showed above in Figure 3. These are the discretized versions of the continuous-time processes given in equations (32) and (33). In these Figures, we also plot the Almgren-Chriss (AC) strategy from Almgren and Chriss, (2001), which is the same benchmark used by Cartea et al., (2015) and part of the motivation for their models. The AC strategy is deterministic and for the acquistion problem it can be defined, as given in Cartea et al., (2015), by the following formula,
@@
Lastly, in Figures 7 and 8, we show heat maps for the inventory and trading speed processes over all 10,000 simulations. The blue line in both graphs shows the mean of all the inventory and trading speed paths under our optimal strategy, while the black line shows the mean of the inventory and trading speed paths under the AC strategy, which recall is a benchmark strategy. In Figure 7, we notice that as we increment either of the parameters $\bar{\sigma}$ and $\varsigma$, the optimal inventory line becomes more concave shaped and thus moves away from the AC line, with this effect increasing the further we are from $t=0$ and $t=T$. In Figure 8, one will notice that as we increment either of the parameters $\bar{\sigma}$ and $\varsigma$, the trading speed will increase significantly earlier on in the simulation, and thus decreases later on as it acquired most of its targeted inventory a lot earlier. In both plots, we can see that there are a lot of points at the top of inventory plot (Figure 7) and the bottom of the trading speed plot (Figure 8). These are the simulations where the barrier at $S_{max}$ or $y=0$ got breached before $t=T$. There are more of these instances as we increment either of the parameter values $\bar{\sigma}$ and $\varsigma$, which would make sense intuitively, since when the overall level of volatility is higher, the agent is induced to complete the acquisition strategy a lot quicker.
@@ 5.2 Liquidation Problem Strategy Simulations
Here, as in section 5.1, we analyse how the strategy performed in our simulations but for the liquidation problem. First, see below, in Figure 9, the first five sample price paths from our simulations and, in Figure 10, a histogram showing the frequency of the average traded price in each of the $10,000$ simulations. In Figure 9, these 5 random price paths are again based on our price process defined earlier in equation (17), where we repeat the simulations for different values of either $\bar{\sigma}$ or $\varsigma$. The dashed black line in both plots now shows the price floor, $S_{min}$, which if breached ends the simulation. As in section 5.1, Figure 10 portrays how the price paths end up influencing the average asset prices the agent ends up liquidating units of the asset at. It is, again, quite obvious that as we increment either $\bar{\sigma}$ or $\varsigma$, prices are more likely to exhibit higher volatility, as this is still a diffusion model, but we will see next how this higher volatility coming from the jump part (via diffusion approximation) can significantly alter how the simulated trading strategy will evolve. One thing to notice from these plots is that when our boundary condition $h(t,S_{min}=19.8)$ is breached. Here, we encounter one of the times where the agent liquidates the remaining units instantly, pays the terminal penalty $\alpha=0.01$ and stops trading. The other time this would occur at is when $t=T=1$, like in the acquisition problem. We will notice in the next set of plots how the strategy is then significantly altered each time this occurs and we will portray how this can significantly effect the overall performance of the strategy in these scenarios.
@@
Next, in Figures 11 and 12, we show how the agents inventory and trading speed processes evolve over the five sample price paths we showed above in Figure 9. These are again the discretized versions of the continuous-time processes given in equations (46) and (47). In these Figures, we again plot the Almgren-Chriss (AC) strategy, where the formula is slightly different to equation (55), as we are now dealing with the liquidation problem. This formula, as defined in Cartea et al., (2015), is,
@@
The common theme, throughout these subplots is that as we increment either $\bar{\sigma}$ or $\varsigma$, these processes themselves become a lot more volatile and are heavily determined by the direction and the degree to which the price is moving along each path. We notice in the top left subplot, that the trading strategy (viewed through the lens of the inventory and trading speed processes) does not vary too much along those different price paths. But as we again introduce nonzero increasing values for either of the coefficients $\bar{\sigma}$ or $\varsigma$ from our models, a large portion of the simulations end up liquidating units of the asset a lot quicker and many have liquidated the targeted inventory well before the maturity time, $T$. This is not too surprising, as the risk of obtaining good liquidation prices decreases as these coefficients increase, thus the agent is induced to trade a lot quicker as a result of this added risk. This is quite important because if a trader were to run a liquidation strategy like this in live markets, and the price were to suddenly jump down, they would want the algorithm to react accordingly. In other words, if a sudden jump down were to enter the market, $\varsigma$ would instantly increase, inducing the trader to increase the speed of liquidating units before the price floor is breached. This increase in the speed of trading would be higher the closer the agent is to the maturity time, $T$. Remember, that if prices were to jump up, this would actually be beneficial to the trader, as they would then end up liquidating at higher prices and there should be no increase in urgency to liquidate units of the asset quickly. Thus, we don’t focus much on this case as the optimal trading speed is then mostly based on the time left to maturity, which would be quite similar to the AC model for the same number of remaining units of the asset left to liquidate.
@@
Lastly, in Figures 13 and 14, we show heat maps for the inventory and trading speed processes over all 10,000 simulations. Like in Figure 7 and 8 for the acquisition strategy simulations, the blue line in both graphs shows the mean of all the inventory and trading speed paths under our optimal strategy, while the black line shows the mean of the inventory and trading speed paths under the AC strategy, which recall is a benchmark strategy. In Figure 13, we notice that as we increment either of the parameters $\bar{\sigma}$ and $\varsigma$, the optimal inventory line becomes more convex shaped and thus moves away from the AC line, with this effect increasing the further we are from $t=0$ and $t=T$. In Figure 14, which is very similar to Figure 8 for the acquisition problem simulations, one will notice that as we increment either of the parameters $\bar{\sigma}$ and $\varsigma$, the trading speed will increase significantly earlier on in the simulation, and thus decreases later on as it liquidated most of its targeted inventory a lot earlier. In both plots, we can see that there are a lot of points at the bottom. These are the simulations where the barrier at $S_{min}$ or $q=0$ got breached before $t=T$. There are more of these instances as we increment either of the parameter values $\bar{\sigma}$ and $\varsigma$. This makes sense intuitively, since when the overall level of volatility is higher, just like in the acquisition problem, the agent is induced to complete the liquidation strategy a lot quicker.

### v2

@@ 4.2 Liquidation Problem Solution
For the liquidation problem, as for the acquisition problem in Section 4.1, we must ensure that we have a well-posed problem. To achieve this, we now specify a boundary condition along $S_{max}\gg S_{min}$, similarly to how a boundary condition was specified along $S_{min}\ll S_{max}$ in Section 4.1. Here, we specify the boundary condition $\partial_{ss}h|_{S=S_{max}}=0$. Then, the numerical scheme, again for $i=1,2,...,M-1$ and $n=0,1,...,N-1$, for the liquidation problem, where its PDE was defined in Equation (43), can be derived as follows:
@@
where $M_{1}$ is defined as in Section 4.1. Note that here one must also account for the boundary conditions at $i=0$, $i=M$ and for the terminal condition at $n=N$. These are as defined in Equations (44)-(45) and the condition above where an upper boundary condition was set at $S_{max}$. One can see that Equation (54) is very similar to Equation (51), which was used to solve the acquisition problem. The main changes here are in the boundary conditions. The liquidation problem now has a price floor rather than a price cap, as in the acquisition problem.
@@
Next, in Figure 3, we show a set of subplots which display contour plots portraying how the solution changes as a function of time and the financial asset’s price, over different values of $\bar{\sigma}$ and $\varsigma$. Time, here, is again from $t=0$ to the maturity time $t=T=1$, and the midprice (the space variable) is from $S_{min}=30.8$ to $S_{0}=30.97$. We specifically focus on this part of the solution as this is where it varies the most. Right of each subplot shows the color scheme for the values of the solution that are displayed. These solution values vary from 00 to $0.01$, where $\alpha=0.01$ represents the value of the terminal condition and the boundary condition at $T$ and $S_{min}$, respectively, as defined in Equations (44)-(45). We can see that the agents speed of liquidating units increases as we get closer to the terminal time. This is because the agent is trying to avoid the terminal penalty cost that they must pay to liquidate any remaining units at maturity. Furthermore, the speed of trading also increases as the asset’s price approaches the floor price, $S_{min}$, again to avoid the other terminal penalty. This trading process was formalized in the stopping time in Equation (35).
@@
As in Figure 2, the format for the pricing models in these plots is similar. In the top left subplot of Figure 3, we show the solution to the liquidation problem for the case where both $\bar{\sigma}=0$ and $\varsigma=0$, which would represent the case for the pure diffusion benchmark models in Cartea et al., (2015). In the top right subplot, we show the solution for the case where $\bar{\sigma}=0.01598$ and $\varsigma=0$, which would portray the same solution were the pricing model to follow the price process in Cartea et al., (2015) combined with the price processes developed in Roldan Contreras and Swishchuk, (2022) and Roldan Contreras, (2023). The two bottom subplots, where we increased the value of the parameter $\varsigma$, where the bold value represents the calibrated parameter value, then show the solutions for the cases where the price process follows our more general jump-diffusion model given in Equation (17).
@@
In each consecutive subplot, from left to right, we again vary either $\bar{\sigma}$ or $\varsigma$ to show the effect that the Semi-Markov and Hawkes process parts can have on the optimal solution. In each of these subplots, one can see the effect on the solution matrix as we slowly incremented the effect of either of the parameters, $\bar{\sigma}$ or $\varsigma$. We can see that as they increase, the speed of trading increases at earlier time steps at lower asset prices. This makes sense intuitively, as the higher these coefficients become, the more overall volatility the solution predicts. Thus, in these scenarios, it would then be more optimal to liquidate units of the asset quicker based on our model. In this particular type of trading problem, where the agent is liquidating units, it’s mainly price jumps that go down which can significantly change the optimal solution as the agent will want to liquidate units of the asset a lot quicker then. In the case where the asset price would jump up, this would be highly favorable to the agent, and the optimal speed of liquidating any remaining units of the asset would then mostly depend on time.
@@
Most of the input parameters for this model are kept the same as in Table 1 in Section 4.1 for the acquisition problem. The only value that we changed was $S_{min}$ and $S_{max}$, which here were set to $30.8$ and $33$, respectively. These were changed so that we could focus more on the impact of the price floor on the optimal solution, whereas in Section 4.1 the solution is more focused on the effect of the price cap.
@@ 5 Strategy Simulations: Acquisition and Liquidation
In this Section, we simulate the performance of the above acquisition and liquidation problems over $10,000$ different price paths, which is the total number of simulations we ran. In Section 4, the optimal solution takes a static view of the markets, whereas in reality, market conditions are constantly evolving. This should be reflected in the agent’s strategy and is throughout our strategy simulations, where the strategy is updated at each time step to reflect this. Here, we will, again, specifically focus on how an increasing $\bar{\sigma}$ or $\varsigma$ can significantly change how the acquisition/liquidation strategy would evolve and how the certain processes we defined in Section 3 are affected. To setup these simulations, one must discretize the continuous-time processes introduced in Section 3. In other words, the continuous-time processes in Equations (14), (17), (18), (19), (21), (32)-(34) and (46)-(48) are discretized. Since both the acquisition and liquidation trading problems have already been solved numerically, as shown in Section 4, this part is already given to us in a discrete form, and we can use the values in these solution matrices to guide our trading decisions in the strategy simulation. As we increment forward in time discretely, each simulation looks at the current price at every time step and then selects the matching value from the $h$ solution matrix, which guides the agent on what the new optimal trading speed is. After a trade is made, the cash, inventory, and execution price processes are updated. If either the terminal or boundary condition is breached, the simulation ends and the agent pays the terminal penalty to acquire/liquidate any remaining units of the asset. In this Section we will again present subplots in the same format as in Figure 1 and 2. Recall that in those Figures, that the top left subplot is for the Cartea et al., (2015) case, the top right subplot is for the Cartea et al., (2015) case combined with the cases in Roldan Contreras and Swishchuk, (2022) and Roldan Contreras, (2023), and the bottom two subplots are for our new more general price processes as defined in Equation (17).
@@ 5.1 Acquisition Problem Strategy Simulations
In this Section we will discuss the strategy simulations for the acquisition problem. These are very similar to the simulations performed in Cartea et al., (2015) in their pure diffusion models. First, see below, in Figure 4, the first five sample price paths from our simulations and, in Figure 5, a histogram showing the frequency of the average traded price in each of the $10,000$ simulations. In Figure 4, these 5 random price paths are based on our price process defined earlier in Equation (17), where we repeat the simulations for different values of $\bar{\sigma}$ or $\varsigma$. In Figure 5, one can see how the different parameter values can influence the asset price the agent ends up acquiring units at in our strategy simulations. The dashed black line in both Figures also shows the price cap, $S_{max}$, which if breached ends the trading strategy simulation.
It is quite obvious that as the $\bar{\sigma}$ or $\varsigma$ values are increment, prices are more likely to exhibit higher volatility, as this is still a diffusion model, but we will see next how this higher volatility coming from the jump part (via diffusion approximation) can significantly alter how the simulated trading strategy will evolve. One thing to notice from these plots is that when our boundary condition $h(t,S_{max}=31.1)$ is breached. Here, we encounter one of the times where the agent purchases the remaining units instantly, pays the terminal penalty $\alpha=0.01$ and stops trading. The other time this can occur is when $t=T=1$. We will notice in the next set of plots how the strategy is then significantly altered each time this occurs and we will portray how this can significantly effect the overall performance of the strategy in these scenarios.
Next, in Figures 6 and 7, we show how the agent’s inventory and trading speed processes evolve over the five sample price paths we showed above in Figure 4. These are the discretized versions of the continuous-time processes given in Equations (32) and (33). In these Figures, we also plot the Almgren-Chriss (AC) strategy from Almgren and Chriss, (2001), which is the same benchmark used by Cartea et al., (2015) and part of the motivation for their models. The AC strategy is deterministic and for the acquistion problem it can be defined, as given in Cartea et al., (2015), by the following formula,
@@
Lastly, in Figures 8 and 9, we show heat maps for the inventory and trading speed processes over all 10,000 simulations. The blue line in both graphs shows the mean of all the inventory and trading speed paths under our optimal strategy, while the black line shows the mean of the inventory and trading speed paths under the AC strategy, which recall is a benchmark strategy. In Figure 8, we notice that as we increment either of the parameters $\bar{\sigma}$ and $\varsigma$, the optimal inventory line becomes more concave shaped and thus moves away from the AC line, with this effect increasing the further we are from $t=0$ and $t=T$. In Figure 9, one will notice that as we increment either of the parameters $\bar{\sigma}$ and $\varsigma$, the trading speed will increase significantly earlier on in the simulation, and thus decreases later on as it acquired most of its targeted inventory a lot earlier. In both plots, we can see that there are a lot of points at the top of inventory plot (Figure 8) and the bottom of the trading speed plot (Figure 9). These are the simulations where the barrier at $S_{max}$ or $y=0$ got breached before $t=T$. There are more of these instances as we increment either of the parameter values $\bar{\sigma}$ and $\varsigma$, which would make sense intuitively, since when the overall level of volatility is higher, the agent is induced to complete the acquisition strategy a lot quicker.
@@ 5.2 Liquidation Problem Strategy Simulations
Here, as in Section 5.1, we analyze how the strategy performed in our simulations but for the liquidation problem. First, see below, in Figure 10, the first five sample price paths from our simulations and, in Figure 11, a histogram showing the frequency of the average traded price in each of the $10,000$ simulations. In Figure 10, these 5 random price paths are again based on our price process defined earlier in Equation (17), where we repeat the simulations for different values of either $\bar{\sigma}$ or $\varsigma$. The dashed black line in both plots now shows the price floor, $S_{min}$, which if breached ends the simulation. As in Section 5.1, Figure 11 portrays how the price paths end up influencing the average asset prices the agent ends up liquidating units of the asset at. It is, again, quite obvious that as we increment either $\bar{\sigma}$ or $\varsigma$, prices are more likely to exhibit higher volatility, as this is still a diffusion model, but we will see next how this higher volatility coming from the jump part (via diffusion approximation) can significantly alter how the simulated trading strategy will evolve. One thing to notice from these plots is that when our boundary condition $h(t,S_{min}=19.8)$ is breached. Here, we encounter one of the times where the agent liquidates the remaining units instantly, pays the terminal penalty $\alpha=0.01$ and stops trading. The other time this would occur at is when $t=T=1$, like in the acquisition problem. We will notice in the next set of plots how the strategy is then significantly altered each time this occurs and we will portray how this can significantly effect the overall performance of the strategy in these scenarios.
@@
Next, in Figures 12 and 13, we show how the agents inventory and trading speed processes evolve over the five sample price paths we showed above in Figure 10. These are again the discretized versions of the continuous-time processes given in Equations (46) and (47). In these Figures, we again plot the Almgren-Chriss (AC) strategy, where the formula is slightly different to Equation (55), as we are now dealing with the liquidation problem. This formula, as defined in Cartea et al., (2015), is,
@@
The common theme throughout these subplots is that as we increment either $\bar{\sigma}$ or $\varsigma$, these processes themselves become increasingly more volatile and are heavily determined by the direction and the degree to which the price is moving along each path. We notice in the top left subplot, that the trading strategy (viewed through the lens of the inventory and trading speed processes) does not vary too much along those different price paths. But as we again introduce nonzero increasing values for either of the coefficients $\bar{\sigma}$ or $\varsigma$ from our models, a large portion of the simulations end up liquidating units of the asset a lot faster and many have liquidated the targeted inventory well before the maturity time, $T$. This is not too surprising, as the risk of obtaining good liquidation prices decreases as these coefficients increase, thus the agent is induced to trade a lot faster as a result of this added risk. This is quite important because if a trader were to run a liquidation strategy like this in live markets, and the price were to suddenly jump down, they would want the algorithm to react accordingly. In other words, if a sudden jump down were to enter the market, $\varsigma$ would instantly increase, inducing the trader to increase the speed of liquidating units before the price floor is breached. This increase in the speed of trading would be higher the closer the agent is to the maturity time, $T$. Remember that if prices were to jump up, this would actually be beneficial to the trader, as they would then end up liquidating at higher prices and there should be no increase in urgency to liquidate units of the asset quickly. Thus, we don’t focus much on this case as the optimal trading speed is then mostly based on the time left to maturity, which would be quite similar to the AC model for the same number of remaining units of the asset left to liquidate.
@@
Lastly, in Figures 14 and 15, we show heat maps for the inventory and trading speed processes over all 10,000 simulations. Like in Figure 8 and 9 for the acquisition strategy simulations, the blue line in both graphs shows the mean of all the inventory and trading speed paths under our optimal strategy, while the black line shows the mean of the inventory and trading speed paths under the AC strategy, which recall is a benchmark strategy. In Figure 14, we notice that as we increment either of the parameters $\bar{\sigma}$ and $\varsigma$, the optimal inventory line becomes more convex shaped and thus moves away from the AC line, with this effect increasing the further we are from $t=0$ and $t=T$. In Figure 15, which is very similar to Figure 9 for the acquisition problem simulations, one will notice that as we increment either of the parameters $\bar{\sigma}$ and $\varsigma$, the trading speed will increase significantly earlier on in the simulation, and thus decreases later on as it liquidated most of its targeted inventory a lot earlier. In both plots, we can see that there are a lot of points at the bottom. These are the simulations where the barrier at $S_{min}$ or $q=0$ got breached before $t=T$. There are more of these instances as we increment either of the parameter values $\bar{\sigma}$ and $\varsigma$. This makes sense intuitively, since when the overall level of volatility is higher, just like in the acquisition problem, the agent is induced to complete the liquidation strategy a lot faster.